Circular No. 6823

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1630–47 S. Dieters, University of Alabama in Huntsville; T. Belloni, Astro-

nomical Institute ‘Anton Pannekoek’, University of Amsterdam; E. Kuulkers, Oxford University; A. Harmon, Marshall Space Flight Center, NASA; and R. Sood, Australian Defence Force Academy, University of New South Wales, report: “We have analyzed RXTE/PCA observations of the present outburst of the x-ray transient 4U 1630–47 (IAUC 6822) taken on Feb. 9.07 and 9.7 UT. We report the discovery of two strong quasiperiodic oscillation (QPO) peaks and band-limited noise in the power spectrum of the 2–14keV x-ray flux. During our observations, the source flux (2–60 keV) was ∼ 60 and ∼ 140 mCrab, respectively. The first QPO peak is asymmetric and has centroid frequencies of 2.67 ± 0.01 and 3.19 ± 0.01 Hz, FWHM of 0.19 ± 0.02 and 0.33 ± 0.02 Hz, and fractional rms amplitudes of 16 ± 4 and 14 ± 3 percent, during the first and second observations, respectively. The second QPO peak has centroid frequencies of 5.59 ± 0.07 and 6.53 ± 0.07 Hz, FWHM of 1.2 ± 0.2 and 1.3 ± 0.3 Hz, and fractional rms amplitudes of 5.5 ± 2 and 6 ± 2 percent, during the first and second observations, respectively. The second peak is most likely a harmonic of the first. The band-limited noise component has a break frequency between 2 and 5 Hz and a fractional rms amplitude of ∼ 30 percent. Our results indicate that 4U 1630–47 is in a rare state (so-called very high state) seen previously only in two other blackhole x-ray transients (GX 339–4 and GS 1124–68).”

SUPERNOVAE Unfiltered CCD magnitudes using GSC reference-star magnitudes by

S. Moretti and S. Tomaselli, Forlí, Italy (0.35-m telescope; MT) and by L. Boschini, Monte di Nese, Bergamo, Italy (0.25-m telescope; B): SN 1997dq in NGC 3810, Feb. 5.96 UT, 16.9 (MT); SN 1997ef in UGC 4107, Feb. 5.94, 17.6 (MT). SN 1997eg in NGC 5012, Feb. 7.15, 15.1 (B). SN 1998A in IC 2627, Feb. 7.01, 16.7 (B). SN 1998C in UGC 3825, Feb. 6.86, 17.1 (B). SN 1998D in NGC 5440, Feb. 7.17, 15.8 (B).

COMET 78P/GEHRELS 2 Visual m_{1} estimates (L = reflector): Jan. 2.24 UT, 11.7 (R. Keen, Mt.

Thorodin, CO, 0.32-m L); 21.98, 11.4 (A. Pereira, Cabo da Roca, Portugal, 0.25-m L); 27.83, 11.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m L); Feb. 1.94, 12.8 (R. J. Bouma, Groningen, The Netherlands, 0.25-m L).

1998 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green